OpenStudy (anonymous):

An object is blasted upward at an initial velocity, Vo, of 240 ft/s. The height, h(t), of the object is a function of time, t (in seconds), and is given by the formula h(t)=V0T-16t^2 . How long will it take the object to hit the ground after takeoff? 16 seconds 15 seconds 7.5 seconds 4 seconds

myininaya (myininaya):

h(16)=240(16)-16(16)^2 h(15)=240(15)-16(15)^2 and so on....

myininaya (myininaya):

im wrong solve the following for t 240t-16t^2+16=0 240t-16t^2+15=0 240t-16t^2+7.5=0 240t-16t^2+4=0

OpenStudy (anonymous):

so what do i do with those numbers?

myininaya (myininaya):

use the quadratic formula or try to factor

OpenStudy (anonymous):

im confused

myininaya (myininaya):

for example: 240t-16t^2+16=0 -16t^2+240t+16=0 t^2-15t-1=0 a=1 b=-15 c=-1 t=[15-sqrt{15^2-4(1)(-1)}]/(2)